# Yoga for Fourier–Mukai partnership

### v2

@@ 3. Preservation of adjointness
Hence, the desired claim regarding bounded cohomology follows. Consequently, we have $\Phi_{K}(D^{b}_{\operatorname{coh}}(Y_{1}))\subseteq D^{b}_{\operatorname{coh}}(Y_{2})$, which implies $K$ is relatively perfect over $Y_{1}$. To show the case of perfect complexes being preserved by integral transforms, we can argue as above and use Lemma˜2.1. ∎
Corollary 3.6.
Consider ˜3.1 where $t$ is an affine faithfully flat morphism. Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) if, and only if, $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
@@
This follows from Lemma˜3.3 and Proposition˜3.5
@@
This is well-known to some, but we include it for convenience. By [Sta25, Tag 0B6S & Tag 0E2Z], we know that $(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$ and $f^{!}\mathcal{O}_{S}$ are respectively relative dualizing complexes for $f^{\prime}$ and $f$. Moreover, [Sta25, Tag 0E2Y] tells us that $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$. However, [Sta25, Tag 0E2W] implies relative dualizing complexes are unique up to isomorphism, so $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}\cong(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$. ∎
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension (e.g. $t=1_{S}$). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In this case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension. Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). So, we have a string of isomorphisms
@@
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ ([GW23, Proposition 22.70])
@@
in $D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$. By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$. Then [DLM25, Lemma 3.2] implies
@@
The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect (see see Corollary˜3.6 and the discussion in Section˜2.2), which completes the proof. ∎
@@
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated (i.e. that intersection of two affine open subschemes is affine, see e.g. [AJPV08, §2]). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. Then $\Phi_{K^{\prime}}$ restricts to $D^{b}_{\operatorname{coh}}$.
@@
Let $U_{i}$ be an affine open cover of $S$. After base changing along each open immersion $U_{i}\to S$, we can form open covers of each $Y_{i}$. To check $\Phi_{K^{\prime}}(E)\in D^{b}_{\operatorname{coh}}(Y_{2})$, it suffices to show the restriction of $\Phi_{K^{\prime}}(E)$ to each component of the induced open cover on $Y_{2}$ has bounded and coherent cohomology. Hence, using Proposition˜3.8 and Lemma˜3.3, we can reduce to the case $S$ is affine. In this setting, the desired claim follows from [DLM25, Lemma 3.1]. Specifically, $\Phi_{K^{\prime}}\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D_{\operatorname{qc}}(Y_{1})$ agrees with the unique functor $\Phi\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$ of loc. cit. (indeed, [Nee21b, Example 0.7] gives a triangulated equivalence between the category of finite cohomological functors on perfect complexes and the bounded derived category of bounded pseudocoherent complexes). ∎
Remark 3.10.
In Corollary˜3.9, we require $S$ to be semiseparated solely so that the open immersion from an affine scheme to $S$ is an affine morphism. This allows us to leverage Proposition˜3.8. Note that semiseparated is a mild constraint because any separated scheme is such.
@@
This is baked into the proof of [DLM25, Proposition 4.3], but we spell it out for convenience. In short, the desired claim follows from the string of natural isomorphisms for all $E\in D_{\operatorname{qc}}(Y_{2})$ and $G\in D_{\operatorname{qc}}(Y_{1})$:
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$. In such a case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. By Lemma˜3.3, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$. Moreover, from Lemma˜3.11, the kernel of the integral transform which is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is given $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$. So, it suffices to show there is an isomorphism
@@
As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension. Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). Hence, we have a string of isomorphisms
@@
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$ ([GW23, Proposition 22.70])
@@
in $D_{\operatorname{coh}}^{+}$. This completes the proof. ∎
@@ 4. Fully faithfulness & equivalences
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\alpha^{K}\colon\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$.

### v3

@@ 3. Preservation of adjointness
Hence, the desired claim regarding bounded cohomology follows. Consequently, we have $\Phi_{K}(D^{b}_{\operatorname{coh}}(Y_{1}))\subseteq D^{b}_{\operatorname{coh}}(Y_{2})$, which implies $K$ is relatively perfect over $Y_{1}$.
Next, we check the second claim. Choose any $p\in Y_{1}$. Denote by $t\colon\operatorname{Spec}(\kappa(f_{1}(p)))\to S$ for the natural morphism. There is a commutative diagram
[…]
where $h_{1}$ is the natural morphism. Using the hypothesis, it follows that
$\mathbf{R}(g^{\prime}_{2})_{\ast}$ $(\mathbf{L}(t^{\prime})^{\ast}\big(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2})\big))$
$\subseteq\mathbf{R}(g^{\prime}_{2})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))))$
$\subseteq\operatorname{Perf}(Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))).$
There is a fibered square
$Y_{1}\times_{S}Y_{2}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))$$Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))$$Y_{1}\times_{S}Y_{2}$$Y_{1}.$$\scriptstyle{g_{2}^{\prime}}$$\scriptstyle{t^{\prime}}$$\scriptstyle{t_{1}}$$\scriptstyle{f^{\prime}_{2}}$
So, from flat base change, we have for all $P\in\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$,
$\mathbf{L}t_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(K\otimes^{\mathbf{L}}P)$
$\in\operatorname{Perf}(Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))).$
This implies that
$\mathbf{L}h_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\cong\mathbf{L}h^{\ast}\mathbf{L}t_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\in D^{b}_{\operatorname{coh}}(\kappa(p)),$
because the derived pullback of perfect complexes remain perfect (see e.g. [STA25, Tag 09UA]). Using [AJS23, Theorem 2.3(iv)], we deduce that $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\in\operatorname{Perf}(Y_{1})$; note that $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ is bounded because $K\otimes^{\mathbf{L}}P$ is a bounded pseudocoherent complex and $f^{\prime}_{2}$ is proper. Indeed, we have shown that $\mathbf{L}b^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ is bounded for all $q\in Y_{1}$ with natural morphism $b\colon\operatorname{Spec}(\kappa(q))\to Y_{1}$. Thus, $K$ is relatively perfect over $Y_{1}$.
Lastly, we check the third claim. However, we can argue essentially like above. In particular, the argument above shows that the stalks of $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ at all closed points of $Y_{1}$ are perfect. Consequently, the desired claim follows from Proposition˜2.4. ∎
Now, for the main result of the section.
Theorem 3.6.
Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then the following are equivalent:
  1. (1)
    $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$)
  2. (2)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for every affine morphism $t$ from a Noetherian scheme
  3. (3)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $t\colon\operatorname{Spec}(k)\to S$ from a field
  4. (4)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for all closed immersions $t\colon\operatorname{Spec}(k(p))\to S$ associated to a closed point $p\in S$
  5. (5)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for some affine surjective morphism $t\colon T\to S$.
@@
We only prove the case of being relatively perfect over $Y_{1}$ because the other can be shown analogously. By Lemma˜3.3, we see that (1) implies (2). Moreover, we know that (2) implies (3) by [STA25, Tag 01SI]. Clearly, (3) implies (4). Furthermore, to see (4) implies (5), use that Item˜3.5(3) tells us $K$ is relatively perfect over $Y_{1}$, i.e. take $t=1_{S}$.
So, we need to check (5) implies (1). Let $t\colon T\to S$ be an affine surjective morphism such that $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$. Choose any morphism $s\colon\operatorname{Spec}(k)\to S$ from a field. Consider the fiber product
[…]
By base change, $s^{\prime}$ is affine and $t^{\prime}$ is an affine surjection. To avoid complicated diagrams involving base changes, we need some notation. Specifically, let $K_{\#}$ be the derived pullback of $K$ along the natural morphism $Y_{1}\times_{S}Y_{2}\times_{S}\#\to Y_{1}\times_{S}Y_{2}$ for #\# any scheme appearing in the diagram above. Now, the hypothesis tells us that $K_{T}$ is relatively perfect over $Y_{1}\times_{S}T$. Hence, Lemma˜3.3 implies $K_{T\times_{S}\operatorname{Spec}(k)}$ is relatively perfect over $Y_{1}\times_{S}T\times_{S}\operatorname{Spec}(k)$ because $s^{\prime}$ is affine. However, $t^{\prime}$ is affine and faithfully flat, so Item˜3.5(1) ensures that $K_{\operatorname{Spec}(k)}$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$. However, $s$ was an arbitrary morphism from a field, so Item˜3.5(2) tells us $K$ is relatively perfect over $Y_{1}$ as desired. ∎
Next, we turn to studying adjointness under base change. First, we focus on right adjoints.
@@
This is known but we include it for convenience. By [STA25, Tag 0B6S & Tag 0E2Z], we know that $(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$ and $f^{!}\mathcal{O}_{S}$ are respectively relative dualizing complexes for $f^{\prime}$ and $f$. Moreover, [STA25, Tag 0E2Y] tells us that $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$. However, [STA25, Tag 0E2W] implies relative dualizing complexes are unique up to isomorphism, and so $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}\cong(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$. ∎
@@
Consider ˜3.1 with $t$ affine. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In both cases, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [STA25, Tag 0AU1]). So, we have a string of isomorphisms
@@
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ (Lemma˜2.2)
@@
By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$. Then [DLM25, Lemma 3.2] implies
@@
The last isomorphism comes from the work above, whereas the second isomorphism is [RIZ17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect via Theorem˜3.6, which completes the proof. ∎
@@
Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. Then $\Phi_{K^{\prime}}$ restricts to $D^{b}_{\operatorname{coh}}$.
@@
By Theorem˜3.6, it suffices to check that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ restricts to $D^{b}_{\operatorname{coh}}$ for every $t\colon\operatorname{Spec}(\kappa(p))\to S$ for $p\in S$ a closed point. So, we can reduce to the case $S=\operatorname{Spec}(k)$ for $k$ a field. In this setting, the desired claim follows from [DLM25, Lemma 3.1]. Specifically, $\Phi_{K^{\prime}}\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D_{\operatorname{qc}}(Y_{1})$ agrees with the unique functor $\Phi\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$ of loc. cit. (indeed, [NEE21b, Example 0.7] gives a triangulated equivalence between the category of finite cohomological functors on perfect complexes and the bounded derived category of bounded pseudocoherent complexes). ∎
Next, we focus on left adjoints.
@@
This is baked into the proof of [DLM25, Proposition 4.3] but we spell it out for convenience. In short, the desired claim follows from the string of natural isomorphisms for all $E\in D_{\operatorname{qc}}(Y_{2})$ and $G\in D_{\operatorname{qc}}(Y_{1})$:
@@
Consider ˜3.1 with $t$ affine. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$. In such a case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
That the last claim holds follows from [RIZ17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. By Lemma˜3.3, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$. Moreover, from Lemma˜3.10, the kernel of the integral transform which is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is given by the object $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$. So, it suffices to show there is an isomorphism
@@
As in Lemma˜3.2, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$ because upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [STA25, Tag 0AU1]). Hence, we have a string of isomorphisms
@@
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$ (Lemma˜2.2)
@@
This completes the proof. ∎
@@ 4. Fully faithfulness & equivalences
Lemma 4.1 ([MAC78, IV.7, Exercise 4]).
Consider a diagram of functors and adjoint functors
(4.1) $\mathcal{C}$$\mathcal{C}^{\prime}$$\mathcal{D}$$\mathcal{D}^{\prime}.$$\scriptstyle{H}$$\scriptstyle{F}$$\scriptstyle{F^{\prime}}$$\scriptstyle{G}$$\scriptstyle{K}$$\scriptstyle{G^{\prime}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
Let $\beta\colon F^{\prime}H\Rightarrow KF$ and $\alpha\colon HG\Rightarrow G^{\prime}K$ be a pair of natural transformations. Then the following conditions are equivalent:
  1. (1)
    The following diagram commutes:
[…]
  2. (2)
    The following diagram commutes:
[…]
  3. (3)
    For all objects $c\in\mathcal{C}$, $d\in\mathcal{D}$, the following diagram commutes:
[…]
Proof.
[…]
$f\mapsto Gf\cdot\eta_{c}\mapsto H(Gf\cdot\eta_{c})\mapsto\alpha_{d}\cdot H(Gf\cdot\eta_{c})=\alpha_{d}\cdot HGf\cdot H\eta_{c}.$
Moreover, $f$ is assigned along the left bottom composite as
$f\mapsto Kf\mapsto Kf\cdot\beta_{c}\mapsto G^{\prime}(Kf\cdot\beta_{c})\cdot\eta^{\prime}_{Hc}=G^{\prime}Kf\cdot G^{\prime}\beta_{c}\cdot\eta^{\prime}_{Hc}.$
So, by the naturality of $\alpha$ and (4.1(2)), we have
$G^{\prime}Kf\cdot G^{\prime}\beta_{c}\cdot\eta^{\prime}_{Hc}=G^{\prime}Kf\cdot\alpha_{Fc}\cdot H\eta_{c}=\alpha_{d}\cdot HGf\cdot H\eta_{c}.$
[…]
Definition 4.2.
A (lax) morphism of adjunctions from $F\dashv G$ to $F^{\prime}\dashv G^{\prime}$ is comprised of a pair of functors
$\mathcal{C}$$\mathcal{C}^{\prime}$$\mathcal{D}$$\mathcal{D}^{\prime}$$\scriptstyle{H}$$\scriptstyle{F}$$\scriptstyle{F^{\prime}}$$\scriptstyle{G}$$\scriptstyle{K}$$\scriptstyle{G^{\prime}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
plus a pair of natural transformations $\beta\colon F^{\prime}H\Rightarrow KF$, $\alpha\colon HG\Rightarrow G^{\prime}K$ such that the equivalent conditions in Lemma 4.1 hold. We call $\beta$ (resp. $\alpha$) the left (resp. right) comparison transformation. Furthermore, a strict morphism of adjunctions is just a lax morphism of adjunctions such that $F^{\prime}H=KF$, $HG=G^{\prime}K$, and $\beta$, $\alpha$ are the identity transformation.
Remark 4.3.
In the category theory literature, a ‘morphism of adjunctions’ is usually referred in its strict sense [MAC78, IV.7] and [RIE17, Exercise 4.2.v], but a ‘lax’ one is instead described by saying that $\beta$ and $\alpha$ are mates under Diagram˜4.1. See [CGR14, Sect. 1] for the precise definition of mates. Indeed, the mate $\overline{\beta}$ of $\beta$ is by definition the natural transformation
$\overline{\beta}\colon HG\xrightarrow{\eta^{\prime}_{HG}}G^{\prime}F^{\prime}HG\xrightarrow{G^{\prime}\beta_{G}}G^{\prime}KFG\xrightarrow{G^{\prime}K\varepsilon}G^{\prime}K.$
Therefore, $\overline{\beta}$ will equal $\alpha$ if, and only if, the corresponding adjunct morphisms $F^{\prime}HG\to K$ (with respect to $F^{\prime}\dashv G^{\prime}$) are equal. This is exactly Item˜4.1(1). Dually, one also sees that Item˜4.1(2) amounts to the mate $\overline{\alpha}$ of $\alpha$ being equal to $\beta$.
@@
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then on $D_{\operatorname{qc}}$ there is a natural isomorphism
$\alpha^{K}\colon\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}\to\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$